# On the Aicardi-Juyumaya bracket for tied links

### v1

@@ Abstract.
[…]
@@ 1. Introduction
[…]
@@
[…]
@@
[…]
Acknowledgements
@@ 2.1. Tied links
The reason why these objects are called tied links is that originally, the components belonging to the same partition were connected by ties, visually represented by dashed lines, as shown in Figure 1
@@ 2.2. Aicardi-Juyumaya bracket
[…]
@@
If we have a crossing of type $D_{+,\sim}$, we will denote by $\bar{0}$ and $\bar{1}$, respectively, the smoothings to obtain the diagrams $D_{0}$ and $D_{\infty}$ when applying axiom $(5)$. On the other hand, if we have a crossing of type $D_{+}$, we will denote by 0, 1, and 2, respectively, the smoothings to obtain the diagrams $D_{0}$, $D_{\infty}$, and $D_{-}$ when applying axiom $(6)$. Figure 4 illustrates this notation:
@@
Observe that when we apply a smoothing of type $\bar{0}$, $\bar{1}$, or $2$, the colors of the involved arcs are preserved. However, when smoothings are of type 00 or $1$, the colors of the arcs of the distinguished crossings merge into one, that is, the two associated blocks of the partition are joined. We stablish the arcs in $D_{0}$ and $D_{\infty}$ to inherit the color of lower index, as shown in Figure 4.
@@
Here, $(-A^{2}-A^{-2})^{k-1}$ is the value of the bracket for each Kauffman state $s$, where $k$ is the number of circles in the smoothed state associated to $s$, and $A^{c-2r}$ is the product of the labels of the edges connecting $s$ to the root, with $c$ the number of crossings of $D$ and $r$ the number of 1-smoothings in the state $s$.
@@
[…]
@@ 3.1. Complexity of tied link diagrams
Diagram $D$ shown in Figure 5 has $4$ illegal crossings: two of them are of type $2$ (color $2$ (grey) is below color $1$ (black)) and two of them of type $1$ (involving black arcs). Therefore, the complexity of $D$ is $(6,4)$.
@@
From Remark 2.4, we know that we reduce the complexity of tied link diagrams when using axioms $(5)$ and $(6)$ until reaching the AJ-states. Indeed, we have that once an illegal crossing $c$ in a diagram of complexity $(n,k)$ is smoothed, if it is of type 1, then two diagrams of complexity $(n-1,k-1)$ are obtained. On the other hand, if $c$ is of type 2, a diagram of complexity $(n,k-1)$ and two diagrams of complexity $(n-1,k^{\prime})$ with $k^{\prime}\in\mathbb{N}$ are obtained. In both cases, one obtains diagrams with lower complexity. Observe that the change of colors of some components might cause $k^{\prime}$ to be greater than $k$.
@@ 3.2. Contribution of an $AJ$-state to $\langle\langle D\rangle\rangle$
      Again, we used the induction hypothesis in the second equality, as in $D_{r(2)}$, $D_{r(0)}$ and $D_{r(1)}$ $l$ is an illegar crossing of type 1. We then notice that $s(D_{r(i),l(j)})=s(D_{l(j),r(i)})$ for every $i\in\{0,1,2\}$ and $j\in\{\bar{0},\bar{1}\}$, so $T_{R_{1}}(D_{r(i),l(j)})=T_{R_{2}}(D_{l(j),r(i)})$ by induction hypothesis, therefore $T_{R_{1}}(D)=T_{R_{2}}(D)$.
@@
Thanks to the previous theorem, we know that for every tied link diagram $D$, we can associate a formal sum $T_{R}(D)=\sum_{s\in S}f(s)\cdot s$ which does not depend on the tree $R$, so we will denote it as $T(D)$. Then, we can define a regular isotopy invariant $P$ as follows:
$P(D)=\left\{\begin{array}[]{lll}c^{\gamma(D)-1}(-A^{2}-A^{-2})^{k(D)-\gamma(D)}&&\mbox{if $D$ is an AJ-state}\\ \displaystyle\sum_{s\in S}f(s)P(s)&&\mbox{otherwise},\end{array}\right.$
where $\gamma(D)$ is the number of colors of the AJ-state $D$ and $k(D)$ its number of components.
[…]
@@
    $P(D\widetilde{\sqcup}\bigcirc)$ $=\sum_{s\in S}f(s\widetilde{\sqcup}\bigcirc)P(s\widetilde{\sqcup}\bigcirc)$
@@
[…]
@@
[…]
@@
[…]
{tblr}
width = colspec = Q[200]Q[200]Q[781], row1 = c, cell21 = c, cell22 = c, cell24 = c, cell31 = c, cell32 = c, cell34 = c, cell41 = c, cell42 = c, cell44 = c, cell51 = c, cell52 = c, cell54 = c, cell61 = c, cell62 = c, cell64 = c, hlines, vlines, $D_{1}$ $D{2}$ $\langle\langle D_{1}\rangle\rangle-\langle\langle D_{2}\rangle\rangle$
@@
L11n356{1,0} L11n434{0,0} $-A^{13}-A^{11}c+2A^{9}+3A^{7}c-3A^{3}c-A+\frac{2c}{A}-\frac{1}{A^{3}}-\frac{3c}{A^{5}}+\frac{3c}{A^{9}}+\frac{2}{A^{11}}-\frac{c}{A^{13}}-\frac{1}{A^{15}}$
@@
L10n79{1,1} L10n95{1,0} $-A^{18}-A^{16}c+A^{12}c-2A^{12}-2A^{10}c-A^{8}c+2A^{6}c+A^{6}+2A^{4}c+2A^{4}+A^{2}-c+2+\frac{2c}{A^{2}}+\frac{c}{A^{4}}-\frac{2c}{A^{6}}-\frac{c}{A^{8}}-\frac{2}{A^{8}}-\frac{1}{A^{10}}$
@@
[…]

### v2

@@ Abstract.
Aicardi and Juyumaya generalized the Kauffman bracket of classical links to the setting of tied links by introducing the double bracket $\langle\langle\cdot\rangle\rangle$. However, the analogues of Kauffman’s states and their relationship to this invariant are not immediately clear. We address this question by defining the Aicardi-Juyumaya states, and show that the contribution of each AJ-state to $\langle\langle\cdot\rangle\rangle$ does not depend on the chosen resolution tree. We also present an algorithm to compute the double bracket of a tied link diagram and use it to find pairs of examples of (oriented) tied links sharing the same Homflypt polynomial but different tied Jones polynomial.
@@ 1. Introduction
On the other hand, Kauffman provided in 1987 a combinatorial definition of the Jones polynomial through the so-called Kauffman’s method (see [6, Theorem 2.6]). For this, he defined what is now known as the Kauffman bracket $\langle\cdot\rangle$, which assigns a polynomial to a link diagram and can be defined as follows:
@@
where $\langle\langle\cdot\rangle\rangle$ denotes the Aicardi-Juyumaya bracket, a generalization of the (classical) Kauffman bracket. However, a similar expression to (1.1) for $\langle\langle\cdot\rangle\rangle$ is still unknown.
@@
The structure of the paper is as follows: In Section 2 we present tied links and the Aicardi-Juyumaya bracket $\langle\langle\cdot\rangle\rangle$. In Section 3 we introduce AJ-states and prove the main result. We also present an algorithm allowing to compute the AJ-bracket of a given tied link diagram. In Section 4 we discuss some examples showing the strength of the Aicardi-Juyumaya bracket.
@@ 2.1. Tied links
The reason why these objects are called tied links is that originally, the components belonging to the same partition were connected by ties, visually represented by dashed lines, as shown in Figure 1.
@@ 2.2. Aicardi-Juyumaya bracket
    $\langle\langle\cdot\rangle\rangle$ is invariant under Reidemeister moves R2 y R3.
@@
If we have a crossing of type $D_{+,\sim}$, we will denote by $\bar{0}$ and $\bar{1}$, respectively, the smoothings to obtain the diagrams $D_{0}$ and $D_{\infty}$ when applying axiom $(5)$. On the other hand, if we have a crossing of type $D_{+}$, we will denote by 0, 1, and 2, respectively, the smoothings to obtain the diagrams $_{0}, DD_{\infty}$, and $D_{-}$ when applying axiom $(6)$. Figure 4 illustrates this notation:
@@
Observe that when we apply a smoothing of type $\bar{0}$, $\bar{1}$, or $2$, the colors of the involved arcs are preserved. However, when smoothings are of type 0 or $1$, the colors of the arcs of the distinguished crossings merge into one, that is, the two associated blocks of the partition are joined. We stablish the arcs in $D_{0}$ and $D_{\infty}$ to inherit the color of lower index, as shown in Figure 4.
@@
Here, $(-A^{2}-A^{-2})^{k-1}$ is the value of the bracket for each Kauffman state $s$, where $k$ is the number of circles in the smoothed state associated to $s$, and $A^{c-2r}$ is the product of the labels of the edges connecting $s$ to the root, with $c$ the number of crossings of $D$ and $r$ the number of $\overline{1}$-smoothings in the state $s$.
@@
3. AJ-states and computation of $\langle\langle\cdot\rangle\rangle$
@@ 3.1. Complexity of tied link diagrams
Diagram $D$ shown in Figure 5 has $4$ illegal crossings: two of them are of type $2$ (color $2$ (blue) is below color $1$ (black)) and two of them of type $1$ (involving black arcs). Therefore, the complexity of $D$ is $(6,4)$.
@@
From Remark 2.4, we know that we reduce the complexity of tied link diagrams when using axioms $(5)$ and $(6)$ until reaching the AJ-states. Indeed, we have that once an illegal crossing $c$ in a diagram of complexity $(n,k)$ is smoothed, if it is of type 1, then two diagrams of complexity (n1,k1)(n-1,k-1) are obtained. On the other hand, if $c$ is of type 2, a diagram of complexity (n,k1)(n,k-1) and two diagrams of complexity $(n-1,k^{\prime})$ with $k^{\prime}\in\mathbb{N}$ are obtained. In both cases, one obtains diagrams with lower complexity. Observe that the change of colors of some components might cause $k^{\prime}$ to be greater than $k$.
@@ 3.2. Contribution of an $AJ$-state to $\langle\langle D\rangle\rangle$
      Again, we used the induction hypothesis in the second equality, as in $D_{r(2)}$, $D_{r(0)}$ and $D_{r(1)}$ $l$ is an illegal crossing of type 1. We then notice that $s(D_{r(i),l(j)})=s(D_{l(j),r(i)})$ for every $i\in\{0,1,2\}$ and $j\in\{\bar{0},\bar{1}\}$, so $T_{R_{1}}(D_{r(i),l(j)})=T_{R_{2}}(D_{l(j),r(i)})$ by induction hypothesis, therefore $T_{R_{1}}(D)=T_{R_{2}}(D)$.
@@
Thanks to the previous theorem, we know that for every tied link diagram $D$, we can associate a formal sum
$T_{R}(D)=\sum_{s\in S}f(s)\cdot s,$
which does not depend on the tree $R$, so we will denote it as $T(D)$.
3.3. A regular isotopy invariant
We now use $T(D)$ to define a function $P$, and we show that it coincides with $\langle\langle\cdot\rangle\rangle$, thus proving that $P$ is a regular isotopy invariant.
Definition 3.5.
Let $D$ be a tied link diagram and $s$ an AJ-state of $D$. We define
$P(s)=c^{\gamma(s)-1}(-A^{2}-A^{-2})^{k(s)-\gamma(s)},$
where $\gamma(s)$ denotes the number of colors of the AJ-state $s$, and $k(s)$ its number of components.
For any tied link diagram $D$, we extend the definition of $P$ by setting
$P(D)=\sum_{s\in S}f(s)\,P(s),$
where $S$ denotes the set of AJ-states of $D$.
Proposition 3.6.
The polynomial $P$ coincides with $\langle\langle\cdot\rangle\rangle$, and therefore is a regular isotopy invariant.
Proof.
We will show that $P$ satisfies axioms $(1)$ to $(6)$ of Theorem 2.3 except for (4) since, as mentioned in [2], (4) can be deduced from the others. Let $D$ be a tied link diagram, then we have that for any AJ-state $s$ of $D$, it holds that the smoothings that lead from $D$ to $s$ are the same as those that lead $D\sqcup\bigcirc$ and $D\widetilde{\sqcup}\bigcirc$ to $s$, since their sets of crossings are exactly the same as $D$. From this it follows that $f(s\sqcup\bigcirc)=f(s\widetilde{\sqcup}\bigcirc)=f(s)$. Moreover, if $S$ is the set of AJ-states of $D$, then the AJ-states of $D\sqcup\bigcirc$ and $D\widetilde{\sqcup}\bigcirc$ are, respectively, $\{s\sqcup\bigcirc\mid s\in S\}$ and $\{s\widetilde{\sqcup}\bigcirc\mid s\in S\}$. Now, let us verify the axioms of $\langle\langle\cdot\rangle\rangle$.
@@ 3.3. A regular isotopy invariant
[…]
@@
3.4. Algorithm to compute $\langle\langle\cdot\rangle\rangle$
@@
4. Some key examples and explicit calculation of $\langle\langle\cdot\rangle\rangle$
@@
So we wondered whether the invariant $\langle\langle\cdot\rangle\rangle$ is stronger than the Kauffman bracket. Subsequently, in discussions with one of the authors of the aforementioned work, we were advised to test examples found in [3], where authors studied the strengt of other polynomial invariants for tied links.
The following table shows some pairs of non-equivalent classical links with three components sharing the same writhe (for both members of each pair). These examples cannot be distinguished by their Homflypt polynomial; however, the double bracket $\langle\langle\cdot\rangle\rangle$ distinguish them, and so does the tied Jones polynomial $\mathcal{J}$. For each link, we took the standard diagram appearing in [9].
$D_{1}$ D2D{2} $\langle\langle D_{1}\rangle\rangle-\langle\langle D_{2}\rangle\rangle$
@@
L11n356{1,0} L11n434{0,0} $A^{13}-A^{11}c+2A^{9}+3A^{7}c-3A^{3}c-A+\frac{2c}{A}-\frac{1}{A^{3}}-\frac{3c}{A^{5}}+\frac{3c}{A^{9}}+\frac{2}{A^{11}}-\frac{c}{A^{13}}-\frac{1}{A^{15}}$
@@
L10n79{1,1} L10n95{1,0} $A^{18}-A^{16}c+A^{12}c-2A^{12}-2A^{10}c-A^{8}c+2A^{6}c+A^{6}+2A^{4}c+2A^{4}+A^{2}-c+2+\frac{2c}{A^{2}}+\frac{c}{A^{4}}-\frac{2c}{A^{6}}-\frac{c}{A^{8}}-\frac{2}{A^{8}}-\frac{1}{A^{10}}$
@@
We have tried to find pairs of tied link diagrams with the same $2$-variable Kauffman polynomial and different tied Jones polynomial. To do so, we computed the value of $\langle\langle\cdot\rangle\rangle$ for all pairs of (oriented) link diagrams in [9] with crossing number at most $11$ having the same Kauffman polynomial, considering the finest possible partition (that is, the number of colors equals the number of components). However, the value of $\mathcal{J}$ is the same for both members in each of those pairs. It is an open question to find pairs of links with the same $2$-variable Kauffman polynomial and different value of $\mathcal{J}$.
Acknowledgements